Two pumps operating together can fill a cistern in 4 hours. If one of the pumps working alone can fill the cistern in 6 hours, how long would it take the other pump working alone?
Found 2 solutions by mananth, scott8148:
Answer by mananth(16946)

Let filling the cistern be = 1 job.
Both pumps can do the job in 4 hours.
So they can do 1/4 of the job in 1 hour.
.
one pump does the job in 6 hours .
So it does 1/6th. of the job in 1 hour.
.
Job done by both - job done by one of them = job done by the other in 1 hour.
.
1/4 - 1/6
=3-2 / 12
=1/12.
The other pump does 1/12th job in 1 hour.
.
So it takes 12 hours to do the complete job of filing the cistern alone.

Answer by scott8148(6628)

the 6 hr pump fills 2/3 of the cistern in 4 hr (4/6 = 2/3)

so the other pump fills 1/3 of the cistern in the same 4 hr ___ (1/3)t = 4 ___ t = 12

it would take the other pump 12 hr to fill the cistern working alone